# NeRF-VO: Real-Time Sparse Visual Odometry with Neural Radiance Fields

### 20 Dec 2023 v1

@@ Abstract
We introduce a novel monocular visual odometry (VO) system, NeRF-VO, that integrates learning-based sparse visual odometry for low-latency camera tracking and a neural radiance scene representation for sophisticated dense reconstruction and novel view synthesis. Our system initializes camera poses using sparse visual odometry and obtains view-dependent dense geometry priors from a monocular depth prediction network. We harmonize the scale of poses and dense geometry, treating them as supervisory cues to train a neural implicit scene representation. NeRF-VO demonstrates exceptional performance in both photometric and geometric fidelity of the scene representation by jointly optimizing a sliding window of keyframed poses and the underlying dense geometry, which is accomplished through training the radiance field with volume rendering. We surpass state-of-the-art methods in pose estimation accuracy, novel view synthesis fidelity, and dense reconstruction quality across a variety of synthetic and real-world datasets, while achieving a higher camera tracking frequency and consuming less GPU memory.
1 Introduction
@@
Accurate pose estimation and 3D scene reconstruction of the environment using images are fundamental challenges in 3D computer vision and essential prerequisites for diverse applications in robotics and mixed reality. Neural Radiance Fields (NeRF) [25] have proven to be an excellent scene representation method for novel view synthesis tasks. The original NeRF employs large multi-layer perceptrons (MLPs) to decode 3D coordinates and ray directions into volume density and color, respectively. However, it fails to represent complex and large scenes, particularly those with fine details. Both training and rendering NeRFs with deep MLPs are computationally intensive and time-consuming. Recent works [4, 16, 26] have sought to accelerate NeRF by replacing deep MLPs with more efficient neural representations. For instance, Instant-NGP [26] achieves a substantial speedup by utilizing a hybrid representation that combines trainable multi-resolution hash encodings (MHE) with shared shallow MLPs. These accelerated neural radiance scene representations open the door for real-time critical SLAM/VO techniques.
@@
NeRF-based scene representations enable high-fidelity photometric and geometric reconstruction and provide reasonable estimates for some unobserved regions, while being highly memory efficient. These representations enable the comprehensive utilization of information from raw images. With neural volume rendering, every pixel is leveraged in the scene optimization process. This has potential to push the performance frontier beyond traditional direct [15, 13, 14, 59] and indirect SLAM [10, 18, 3, 6] methods. Lately, numerous works have aimed at integrating SLAM with neural implicit mapping [38, 57, 34, 45, 51, 24]. However, only a few focus on monocular RGB input [32, 58, 21, 5, 55, 54]. In terms of accuracy, RGB-only methods lag behind their RGB-D counterparts, especially in dense reconstruction. Furthermore, most of these approaches, regardless of input modality, are computationally expensive, lack real-time capability, and require significant GPU memory.
@@
To address these issues, we propose NeRF-VO, a real-time capable sparse visual odometry with neural implicit dense mapping. A preview of its performance is provided in Fig. 1. We obtain the initial pose estimation and 3D sparse landmarks using low-latency learning-based sparse visual odometry. Up to scale dense geometric cues, including monocular dense depth and normals, are inferred using a transformer-based neural network. With the initial poses, camera-captured monocular images, and rough dense geometric priors, we can efficiently optimize a neural radiance field that implicitly represents the 3D scene. Accurate poses and dense geometry of the scene are recovered by minimizing the disparity of captured images and predicted dense geometric cues relative to the renderings generated from the neural radiance field. Hence, our proposed NeRF-VO comprises three main components: a sparse visual tracking front-end, a dense geometry enhancement module, and a NeRF-based dense mapping back-end. The system architecture is depicted in Fig. 2.
@@
We present a novel paradigm for optimizing a NeRF scene representation and camera poses by incorporating dense depth and surface normal supervision. Utilizing a transformer-based monocular depth network, we predict dense depth and surface normal priors. To align their scale with sparse visual odometry, we propose a dedicated sparse-to-dense scale alignment procedure.
@@
We open-source our versatile neural SLAM framework, fostering further research by allowing easy integration of other neural representations and VO/SLAM methods.
In the remainder of this paper, we review related work in Sec. 2 and introduce the sparse visual tracking front-end in Sec. 3. We present the main methodology, which includes dense geometry enhancement, scale alignment, neural implicit mapping and dedicated NeRF optimization, in Sec. 4. We report and analyze extensive experimental results in Sec. 5. Finally, we conclude the paper and provide an overview of future work in Sec. 6.
2 Related Work
@@
Learning-Based Visual Tracking. Our work centers on monocular visual odometry using RGB image sequences from a calibrated camera. This technique estimates the camera positions and orientations of each incoming frame. Unlike SLAM, visual odometry focuses on local coherence among consecutive frames and does not include SLAM’s loop closure optimization or global bundle adjustment. Recently, VO methods have advanced from traditional handcrafted feature detection and matching modules to deep learning-based approaches, enhancing accuracy and robustness [40, 50, 46, 29, 41, 42]. Among all of them, DROID-SLAM [41] and DPVO [42] are two noteworthy works that leverage neural networks to predict the optical flow between consecutive images and iteratively update camera poses. DPVO [42] serves as the foundation for the front-end of our SLAM system since it’s highly efficient and accurate.
@@
Dense Visual SLAM. Dense visual SLAM aims to construct a dense 3D representation of the environment instead of sparse 3D landmarks. Following the first real-time dense visual work DTAM [28], many approaches have been proposed, primarily those exploiting monocular depth prediction [7, 56, 44]. The scene representations selected in these works have also progressed from volumetric representations [28] to low-dimension latent representations [2, 7, 60] and the integration of pre-trained depth estimators [40, 19, 49]. In this work, we choose a volumetric neural radiance field as the representation due to its superior photometric and geometric accuracy.
@@
NeRF-Enabled SLAM. Recently, many works have been proposed to integrate NeRF-based [25] neural implicit representations into SLAM. In general, existing methods can be differentiated into one-stage and two-stage approaches. Two-stage approaches use an existing SLAM algorithm as a tracking module to estimate depth maps and camera poses, and then use these estimates as supervisory signals to optimize an implicit neural representation as part of a mapping module. Early approaches such as Orbeez-SLAM [5] and NeRF-SLAM [32] demonstrated the effectiveness of this combination. Recent advancements in this direction have further introduced view-centric implicit functions [23], global loop closure [55, 54], and monocular depth priors [54] to improve pose estimation accuracy and dense mapping quality. In contrast, one-step approaches use a single implicit neural function for both tracking and mapping. iMAP [38] and NICE-SLAM [57] set the groundwork for this direction using RGB-D data. Subsequent works aimed at enhancing the scene representation [34, 45, 51], introducing implicit semantic encoding [24], and integrating inertial measurements [22]. Among these works, NICER-SLAM [58] and DIM-SLAM [21] perform dense SLAM only using monocular images. NICER-SLAM prioritizes high-fidelity scene reconstruction and novel view synthesis through heavy optimization with various losses but is unsuitable for real-time applications. DIM-SLAM, on the other hand, focuses on accurate camera tracking, yet shows suboptimal performance in dense reconstruction.
@@
Our work follows a two-stage design. We use a sparse visual tracking method, DPVO [42] to get initial poses at high frequency and obtain dense depth and surface normals priors from a mono-depth prediction network. These initial poses and dense geometry cues are used to train our mapping back-end employing a nerfacto-based neural radiance field [39]. Our system yields high-frequency pose estimations and promising dense reconstruction after neural scene optimization, all achieved at a low memory footprint.
3 Sparse Visual Pose Tracking
@@
Patch Graph. Given a sequence of RGB frames, DPVO randomly samples a set of square patches of size per keyframe and adds them to a bipartite patch graph that connects patches and frames. For instance, the -th square patch from frame represented by is connected via edges to all frames within temporal vicinity of frame . represent the pixel coordinates and denotes the inverse depths with . is a vector filled with ones. Thus, the patch graph builds a trajectory for each patch, incorporating all of its reprojections , where , and is the array of frames temporally adjacent to . Assuming uniform depth across each patch, the reprojection is defined by:
@@
Differentiable Bundle Adjustment. The key component of DPVO is its differentiable pose and depth update operator. A recurrent neural network operates on the patch graph with a set of edges , while maintaining a hidden state for each edge (patch-frame pair), and predicting a 2D correction vector for each reprojection of the patch center with a corresponding confidence weight . Bundle adjustment is performed using the optical flow correction as a constraint to iteratively update frame poses and patch depths via nonlinear least-squares optimization. The cost function for bundle adjustment is:
where denotes the Mahalanobis distance and denotes the patch center. The bundle adjustment step of DPVO is differentiable, the recurrent neural network is trained on a set of ground truth poses, together with the corresponding ground truth optical flow to supervise pose estimation and optical flow correction.
Keyframing Strategy. DPVO optimizes the patch depths and camera poses of all keyframes in a sliding optimization window, adding each incoming frame. The sliding window keeps a constant array of the most recent keyframes and removes old ones to maintain bounded computational complexity and to ensure runtime efficiency. The three most recent frames are always considered to be keyframes, but the fourth most recent keyframe is only kept if the optical flow to its predecessor is sufficiently high, otherwise, it is removed. We use this keyframe selection for our downstream dense geometry enhancement and mapping modules. To stabilize NeRF optimization, the three most recent keyframes are not considered, as they are likely to be removed later by the keyframing scheme described above. Each time the fourth most recent keyframe is secured, it is propagated to our dense geometry enhancement module and NeRF optimization, including all previously active camera poses and patches in the current sliding window. Since only certain keyframes are involved in our dense geometry enhancement and NeRF optimization, the computational and memory footprint of NeRF-VO are bounded.
4 Neural Implicit Dense Mapping
As indicated in the system overview in Fig. 2, the keyframes from the aforementioned visual tracking front-end are subsequently processed by our dense geometry enhancement and incorporated into our NeRF optimization back-end.
4.1 Dense Geometry Enhancement
@@
Scale Alignment. Since both sparse and dense depths are predicted solely based on RGB input, their depth distributions are skewed relative to the metric depth and to each other. However, downstream dense mapping necessitates depth maps and camera poses at a unified scale. Therefore, we introduce a dense-to-sparse scale alignment procedure to align the predicted dense depth map of each keyframe to the sparse patch depth. Assuming that the sparse depth and the dense depth follow a Gaussian distribution with and , we compute the aligned dense depth by:
@@
where and are the mean and standard deviation statistics of the sparsified dense depth map extracted from at the pixel coordinates of . Since the pixel coordinates are sampled randomly, given a sufficiently large sample size per frame, one could assume that . This would lead to a relaxed formulation of with . We have empirically evaluated this relaxed alignment scheme along with several other alignment strategies and found that the standard formulation in Eq. 4 performs best (see Sec. 5.8).
Outlier Rejection. Due to the randomness in the patch selection, some of the patches are sampled on textureless or non-Lambertian surfaces, causing the bundle adjustment to fail or to produce poor depth estimates. Thus, we introduce an outlier rejection scheme where the patches with depth values in the upper and lower th percentile per frame are excluded from the scale alignment.
@@
4.2 Mapping with Neural Scene Representation
Neural Radiance Field. Our dense mapping back-end reconstructs the scene by optimizing a neural radiance field built on the nerfacto model [39]. The neural scene representation comprises multi-resolution hash encodings similar to [26], spherical harmonics encoding, and two multi-layer perceptrons (MLPs) for density and color recovery. The NeRF scene representation is optimized using the camera poses, captured RGB images as well as inferred depth maps and surface normals of all keyframes. How keyframes are added to this optimization is explained in Sec. 3.
@@ 4.2 Mapping with Neural Scene Representation
The normal supervision is guided by the normal consistency loss proposed in [52], which integrates the L1 distance and cosine distance between the rendered surface normals and the normals predicted from the mono-depth network at pixel :
@@
4.3 System Design
The sparse tracking, dense geometry enhancement, and dense mapping modules run asynchronously on separate threads. Information is propagated in a feed-forward fashion through the three components sequentially. Communication occurs each time a new keyframe is secured in the visual tracking module. It is then processed by the enhancement module and included in the training database of the mapping module. Separated from this process, the mapping module continuously optimizes the implicit representation jointly with the camera poses. With this multi-threaded architecture, our system can run in real-time, with the visual tracking running at low latency, while the enhancement and mapping run at a relatively low frequency, which does not block the data streamed in from the sensor. Due to its asynchronous architecture, our system has the potential to be run on multiple GPUs, leading to improved runtime efficiency.
5 Experiments
We evaluate our method qualitatively and quantitatively against the state-of-the-art monocular SLAM methods using neural implicit representations on a variety of real and synthetic indoor datasets, and present an extensive ablation study of the key design choices of our method.
5.1 Evaluation Datasets

### 16 Jul 2024 v2

@@ Abstract
We introduce a novel monocular visual odometry (VO) system, NeRF-VO, that integrates learning-based sparse visual odometry for low-latency camera tracking and a neural radiance scene representation for fine-detailed dense reconstruction and novel view synthesis. Our system initializes camera poses using sparse visual odometry and obtains view-dependent dense geometry priors from a monocular prediction network. We harmonize the scale of poses and dense geometry, treating them as supervisory cues to train a neural implicit scene representation. NeRF-VO demonstrates exceptional performance in both photometric and geometric fidelity of the scene representation by jointly optimizing a sliding window of keyframed poses and the underlying dense geometry, which is accomplished through training the radiance field with volume rendering. We surpass SOTA methods in pose estimation accuracy, novel view synthesis fidelity, and dense reconstruction quality across a variety of synthetic and real-world datasets while achieving a higher camera tracking frequency and consuming less GPU memory.
Index Terms:
NeRF, Visual Odometry, SLAM, Dense MappingI Introduction
@@
Accurate pose estimation and 3D reconstruction of the environment using images are fundamental challenges in 3D computer vision and essential prerequisites for diverse applications in robotics. Neural Radiance Fields (NeRF) [3] have proven to be an excellent scene representation method for novel view synthesis tasks. The original NeRF employs large multi-layer perceptrons (MLP) to decode 3D coordinates and ray directions into volume density and color. However, it fails to represent complex and large scenes, particularly those with fine detail. Both training and rendering NeRFs with deep MLPs are computationally intensive and time-consuming. Recent works [4, 5, 6] have sought to accelerate NeRFs by replacing deep MLPs with more efficient representations. For instance, Instant-NGP [6] achieves a substantial speedup by utilizing a hybrid representation that combines trainable multi-resolution hash encodings (MHE) with shared shallow MLPs. These accelerated neural implicit scene representations pave the way for real-time critical SLAM/VO.
@@
NeRF-based scene representations enable high-fidelity photometric and geometric reconstructions while being very memory efficient. They enable comprehensive exploitation of information from raw images, leveraging every pixel in the optimization process. This has the potential to push the performance frontier beyond traditional direct [7, 8, 9] and indirect SLAM [10, 11, 12]. Lately, numerous works have aimed at integrating SLAM with neural implicit mapping [2, 13, 14, 15, 16, 17]. Yet, only a few focus on monocular RGB input [18, 19, 20, 21, 22, 23]. In terms of accuracy, RGB-only methods lag behind their RGB-D counterparts, especially in dense reconstruction. Further, most of these approaches are computationally expensive, lack real-time capability, and require significant GPU memory.
@@
To address these issues, we propose a real-time capable sparse visual odometry with neural implicit dense mapping. A preview of its performance is provided in Fig. LABEL:fig:teaser. We obtain the initial pose estimation and sparse 3D landmarks using low-latency learning-based sparse visual odometry. Up-to-scale dense geometric cues, including monocular dense depth and normals, are inferred using a transformer-based neural network. With the initial poses, camera-captured monocular images, and dense geometric priors, we can efficiently optimize a neural radiance field that implicitly represents the 3D scene. Accurate poses and dense geometry of the scene are recovered by minimizing the disparity of captured images and predicted dense geometric cues relative to the renderings generated from the neural radiance field. Hence, our proposed method comprises three main components: a sparse visual tracking front-end, a dense geometry enhancement module, and a NeRF-based dense mapping back-end. The system architecture is depicted in Fig. 1.
@@
We present a novel paradigm for optimizing a NeRF scene representation and camera poses by incorporating dense depth and surface normal supervision. Utilizing a monocular transformer-based network, we predict dense depth and surface normal priors. To align their scale with the sparse visual odometry, we propose a dedicated sparse-to-dense scale alignment procedure.
@@
II Related Work
For brevity, this section covers only the related work most relevant to our approach. Please refer to the survey by Tosi et al. [24] for a more detailed overview of recent advances in NeRF- and GS (Gaussian Splatting)-based SLAM.
@@ II Related Work
Learning-Based Visual Tracking. Our work centers on monocular visual odometry using RGB image sequences from a calibrated camera. This technique estimates the camera positions and orientations of each incoming frame. Unlike SLAM, visual odometry focuses on local coherence among consecutive frames and does not include SLAM’s loop closure optimization or global bundle adjustment. Recently, VO methods have advanced from traditional handcrafted feature detection and matching to deep learning-based approaches, enhancing accuracy and robustness [25, 26, 27, 28, 29]. Among all of them, DROID-SLAM [28] and DPVO [29] are two noteworthy works that leverage neural networks to predict the optical flow between consecutive images and iteratively update camera poses. DPVO [29] serves as the foundation for our system’s front-end since it is highly efficient and accurate.
@@
Dense Visual SLAM. Dense visual SLAM aims to construct a dense 3D representation of the environment instead of sparse 3D landmarks. Following the first real-time dense visual work DTAM [30], many approaches have been proposed, primarily those exploiting monocular depth prediction [31, 32, 33]. The scene representations selected in these works have also progressed from volumetric representations [30] to low-dimension latent representations [34, 31, 35] and the integration of pre-trained depth estimators [25, 36, 37]. In this work, we choose a volumetric NeRF as the representation due to its superior photometric and geometric accuracy.
@@
NeRF-Enabled SLAM. Recently, many works have been proposed to integrate NeRF-based [3] scene representations into SLAM. In general, existing methods can be differentiated into one-stage and two-stage approaches. Two-stage approaches use an existing SLAM algorithm as a tracking module to estimate depth and camera poses, and then use these estimates as supervisory signals to optimize an implicit neural representation as part of a mapping module. Early approaches such as Orbeez-SLAM [21] and NeRF-SLAM [18] demonstrated the effectiveness of this combination. Recent advances in this direction have further introduced view-centric implicit functions [38], global loop closure [22, 23], and monocular depth priors [23] to improve pose estimation accuracy and dense mapping quality. In contrast, one-stage approaches use a single implicit neural function for both tracking and mapping. iMAP [2] and NICE-SLAM [13] set the groundwork for this direction using RGB-D data. Subsequent works aimed at enhancing the scene representation [14, 15, 16], introducing implicit semantic encoding [17], and integrating inertial measurements [39]. Among these, NICER-SLAM [19] and DIM-SLAM [20] perform dense SLAM only using monocular images. NICER-SLAM prioritizes scene reconstruction and novel view synthesis through heavy optimization with various losses, while DIM-SLAM focuses on accurate camera tracking, but shows suboptimal performance in dense reconstruction. However, both methods are not suitable for real-time applications.
GS-Based SLAM. Recently, 3D Gaussian Splatting (3DGS), introduced by Kerbl et al. [40], has attracted attention for its superior ability to model novel view synthesis using differentiable 3D Gaussian-shaped primitives. Recently, this novel paradigm has inspired a variety of SLAM approaches: Analogous to the NeRF-based one-stage approaches, Gaussian Splatting SLAM [41], GS-SLAM [42], and SplaTAM [43] all track the camera using a dense photometric error, while Photo-SLAM [44] uses ORB-SLAM3 [11] for tracking, similar to the NeRF-based two-stage approaches. However, 3DGS does not explicitly model geometry, which hinders its use in robotic tasks that require geometric space awareness.
@@
Our work follows a two-stage design. We use a sparse visual tracking method, DPVO [29] to get initial poses at high frequency and obtain dense depth and surface normals priors from a monocular prediction network. These initial poses and dense geometry cues are used to train our mapping back-end employing a nerfacto-based neural radiance field [45]. Our system yields high-frequency pose estimations and promising dense reconstruction after neural scene optimization, all achieved at a low memory footprint.
III Sparse Visual Pose Tracking
@@
Patch Graph. Given a sequence of RGB frames, DPVO randomly samples a set of square patches of size pixels per keyframe and adds them to a bipartite patch graph that connects patches and frames. For instance, the -th square patch from frame represented by is connected via edges to all frames within temporal vicinity of frame . represent the pixel coordinates and the inverse depths with . is a vector filled with ones. The patch graph builds a trajectory for each patch, incorporating all of its reprojections , where , and is the array of frames temporally adjacent to . Assuming uniform depth across each patch, the reprojection is defined by:
@@
Differentiable Bundle Adjustment. The key component of DPVO is its differentiable pose and depth update operator. A recurrent neural network operates on the patch graph with a set of edges , while maintaining a hidden state for each edge (patch-frame pair), and predicting a 2D correction vector for the reprojection of the patch center with a corresponding confidence weight . Bundle adjustment is performed using the optical flow correction as a constraint to iteratively update frame poses and patch depths via nonlinear least-squares optimization. The cost function for bundle adjustment is:
where denotes the Mahalanobis distance. The bundle adjustment step of DPVO is differentiable, the recurrent neural network is trained on a set of ground truth poses, together with the corresponding ground truth optical flow to supervise pose estimation and optical flow correction.
Keyframing Strategy. DPVO optimizes the keyframes’ patch depths and camera poses in a sliding optimization window, adding each incoming frame. This sliding window keeps a constant array of the most recent keyframes and removes old ones to maintain bounded computational complexity ensuring runtime efficiency. The three most recent frames are always considered to be keyframes, but the fourth most recent keyframe is only kept if the optical flow to its predecessor is sufficiently high, otherwise, it is removed. We use this keyframe selection for our downstream geometry enhancement and mapping modules. To stabilize NeRF optimization, the last three keyframes are not propagated yet, as they are likely to be removed later by the keyframing scheme described above. Each time the fourth most recent keyframe is secured, it is passed to our dense geometry enhancement module and NeRF optimization, including all camera poses and patches in the current sliding window. Since only certain keyframes are involved in our dense geometry enhancement and NeRF optimization, the computational and memory footprint of NeRF-VO are bounded.
IV Neural Implicit Dense Mapping
As indicated in the system overview in Fig. 1, the keyframes from the visual tracking front-end are subsequently processed by our dense geometry enhancement and incorporated into our NeRF optimization back-end.
IV-A Dense Geometry Enhancement
@@
Scale Alignment. Since both sparse and dense depths are predicted solely based on RGB input, their depth distributions are skewed relative to the metric depth and each other. However, downstream dense mapping necessitates depth maps and camera poses at a unified scale. Therefore, we introduce a dense-to-sparse scale alignment procedure to align the predicted dense depth map of each keyframe to the sparse patch depth. Assuming that the sparse depth and the dense depth follow a Gaussian distribution with and , we compute the aligned dense depth by:
@@
where and are the mean and standard deviation of the sparsified dense depth map extracted from at the pixel coordinates of . Since the patches are sampled randomly, given a sufficiently large sample size per frame, one could assume that . This would lead to a relaxed formulation of with . We have empirically evaluated this relaxed alignment scheme along with several other alignment strategies and found that the standard formulation in Eq. 4 performs best (see Sec. V-H).
@@
IV-B Mapping with Neural Scene Representation
Neural Radiance Field. Our dense mapping back-end reconstructs the scene by optimizing a neural radiance field built on the nerfacto model [45]. The neural scene representation comprises multi-resolution hash encodings similar to [6], spherical harmonics encoding, and two multi-layer perceptrons (MLPs) for density and color recovery. The NeRF scene representation is optimized using the camera poses, captured RGB images as well as inferred depth maps and surface normals of all keyframes.
@@ IV-B Mapping with Neural Scene Representation
The normal supervision is guided by the normal consistency loss proposed in [50], which integrates the L1 distance and cosine distance between the rendered surface normals and the normals predicted from the monocular prediction network at pixel :
@@
IV-C System Design
The sparse tracking, dense geometry enhancement, and dense mapping modules run asynchronously on separate threads. Information is propagated in a feed-forward fashion through the three components sequentially. Communication occurs each time a new keyframe is secured in the visual tracking module. It is then processed by the enhancement module and included in the training database of the mapping module. Separated from this process, the mapping module continuously optimizes the implicit representation jointly with the camera poses. With this multi-threaded architecture, our system can run in real-time, with the visual tracking running at low latency, while the enhancement and mapping run at a relatively low frequency, which does not block the data streaming in from the sensor. Due to its asynchronous architecture, our system has the potential to be run on multiple GPUs, leading to improved runtime efficiency.
V Experiments
We evaluate our method qualitatively and quantitatively against the state-of-the-art monocular SLAM methods using neural implicit representations on a variety of real and synthetic indoor datasets and present an extensive ablation study of the key design choices of our method.
V-A Evaluation Datasets